The change in enthalpy (ΔHorxn) for a reaction is -36.6 kJ/mol . The equilibrium constant for the reaction is 2.0×103 at 298 K.

What is the equilibrium constant for the reaction at 621 K ?
Answer #1

ΔHorxn = - 36.6 kJ/mol

temperature T1 = 298 K

K1 = 2.0 x 10^3

T2 = 621 K

ln (K2 / K1) = ΔHo / R [1 / T1 - 1 / T2]

ln (K2 / 2.0 x 10^3) = - 36.6 / 8.314 x 10^-3 [1 / 298 - 1/ 621]

K2 = 0.92

equilibrium constant = 0.92